OpenStudy (anonymous):

Solve for d. -10 = -2 + 1/2 d A. d = -24 B. d = -16 C. d = -6 D. d = -4

OpenStudy (johnweldon1993):

Add 2 to both sides first.... -10 + 2 = -2 + 2 + 1/2d -8 = 1/2d Now multiply 2 to both sides to solve for just 'd' d = -8 * 2 d = ???

OpenStudy (anonymous):

-16 (:

OpenStudy (johnweldon1993):

There you go :)

OpenStudy (anonymous):

\[-10=-2 + \frac{ d }{ 2 }\] So add 2 to both sides \[-8=\frac{ d }{ 2 }\] times both sides by 2 Hope that helps :)